OpenStudy (anonymous):

simplify and evaluate please <3 -7x-3(9-7x) if x=-1 6(3x-1)-10x if x=7 7x-4(6-x)+12 if x=6 -3(4x-1)+9x+8x if x =-9 10-(2y-6)-y if y=8

OpenStudy (tiffanymak1996):

which one do you want as an example?

OpenStudy (anonymous):

well i need help with all since i have no clue what i am doing..

OpenStudy (tiffanymak1996):

well they're kinda the same, you'll just have to substitute the variable in to the equation.

OpenStudy (anonymous):

thats how you simplify ?

OpenStudy (anonymous):

:)

OpenStudy (tiffanymak1996):

that's how to evaluate. when you simplify, :

OpenStudy (tiffanymak1996):

take -7x-3(9-7x) if x=-1 as an example

OpenStudy (tiffanymak1996):

you multiple -3(9-7x) first.

OpenStudy (tiffanymak1996):

can u do that?

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

ya one sec. hi james.. haha

OpenStudy (anonymous):

high

OpenStudy (anonymous):

okay im back so its -3(2x) right??

OpenStudy (anonymous):

?

OpenStudy (anonymous):

http://www.wolframalpha.com

OpenStudy (anonymous):

i dont know what im doing... hahahaa

OpenStudy (tiffanymak1996):

you should multiply it term by term, since 9 and -7x are not like terms.

OpenStudy (anonymous):

what the actual heck

OpenStudy (anonymous):

thank you so much :)

OpenStudy (tiffanymak1996):

-3(9-7x) = -3(9)+(-3)(-7x)= ? can u do the next step?

OpenStudy (anonymous):

no cause i have no clue what the next step is

OpenStudy (tiffanymak1996):

ok, what is -3 x9

OpenStudy (anonymous):

-27

OpenStudy (tiffanymak1996):

yes, what is -3(-7x)?

OpenStudy (anonymous):

21x?

OpenStudy (tiffanymak1996):

yes. so add them up together.

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

or wait

OpenStudy (tiffanymak1996):

go on

OpenStudy (anonymous):

it would be that right? or am i wrong once again

OpenStudy (tiffanymak1996):

remember that -27 and 21x aren't like terms

OpenStudy (anonymous):

you are seriously helping a retard

OpenStudy (tiffanymak1996):

don't feel that way, it should be -27+21x

OpenStudy (tiffanymak1996):

ok, so when -3(9-7x) is -27 +21x when you put that in -7x-3(9-7x), it should become -7x-27+21x, is that ok?

OpenStudy (anonymous):

yea i guess but this is still confusing

OpenStudy (tiffanymak1996):

it's ok, we can do another one later, but now let's finish up this one first.

OpenStudy (anonymous):

okaaayy

OpenStudy (tiffanymak1996):

in -7x-27+21x ; -7x and +21x are like terms, so what is 21x-7x ?

OpenStudy (anonymous):

14x

OpenStudy (tiffanymak1996):

yes, that's right

OpenStudy (tiffanymak1996):

so what is -27+(21x-7x) ?

OpenStudy (anonymous):

-13?

OpenStudy (tiffanymak1996):

you should separate -27 and 14x, since -27 doesn't have a x with it they can't be combined together.

OpenStudy (tiffanymak1996):

so it should be 14x-27 and that's it, do you grasp the idea?

OpenStudy (tiffanymak1996):

when the term has x in it, combine them, when one has a x but the other don't separate them.

OpenStudy (anonymous):

nope i had no clue. lol should i just copy someones homework like i usually do?

OpenStudy (anonymous):

thanks for trying though!

OpenStudy (tiffanymak1996):

don't copy, you can't copy in exams, here: 2x and 5x these are like terms, 2x and 7 these aren't like terms

OpenStudy (anonymous):

i understand that part its just hard i havnt done this algebra crap in likea year.

OpenStudy (tiffanymak1996):

it's ok, you can combine like terms, but you can't combine unlike terms, is this ok?

OpenStudy (anonymous):

yesss mam

OpenStudy (tiffanymak1996):

:) so if 2x+5x=7x

OpenStudy (tiffanymak1996):

and 2x+7 = 2x+7 it stays the same since there isn't any like terms in the question.

OpenStudy (anonymous):

that is easy enough haha

OpenStudy (tiffanymak1996):

yes, so in -27+(21x-7x) are there any like terms and what are they?

OpenStudy (anonymous):

21x and -7x?

OpenStudy (tiffanymak1996):

yes, then you combine them which you get is 14x above, right?

OpenStudy (anonymous):

yessss

OpenStudy (tiffanymak1996):

when -27+(21x-7x) becomes -27 +14x does it have any like terms anymore?

OpenStudy (anonymous):

noooooopppee

OpenStudy (tiffanymak1996):

yup, so the answer stays that way, correct?

OpenStudy (anonymous):

yyeessss mam

OpenStudy (tiffanymak1996):

and that's how you simplify your question, got it?